### v2

# Generating Symbolic Music from Natural Language Prompts using an LLM-Enhanced Dataset

Abstract
Recent years have seen many audio-domain text-to-music generation models that rely on large amounts of text-audio pairs for training. However, symbolic-domain controllable music generation has lagged behind partly due to the lack of a large-scale symbolic music dataset with extensive metadata and captions. In this work, we present MetaScore, a new dataset consisting of 963K musical scores paired with rich metadata, including free-form user-annotated tags, collected from an online music forum. To approach text-to-music generation, we leverage a pretrained large language model (LLM) to generate pseudo natural language captions from the metadata. With the LLM-enhanced MetaScore, we train a text-conditioned music generation model that learns to generate symbolic music from the pseudo captions, allowing control of instruments, genre, composer, complexity and other free-form music descriptors. In addition, we train a tag-conditioned system that supports a predefined set of tags available in MetaScore. Our experimental results show that both the proposed text-to-music and tags-to-music models outperform a baseline text-to-music model in a listening test, while the text-based system offers a more natural interface that allows free-form natural language prompts.
Index Terms:
controllable music generation, multimodal learning, deep neural network, large language modelsI Introduction
Recent work has been investigating the potential of conditional music generation with state-of-the-art machine learning models. In particular, we have seen major progress in audio-domain controllable music generation [24, 23], largely thanks to the vast amount of text-audio pairs for training. Unlike audio-domain music generation, symbolic music generation systems generate music in editable formats that can be further completed by the users, making it easier for musicians to integrate such systems into their creative workflow. However, symbolic-domain controllable music generation has been hindered by the lack of a large, public symbolic music dataset with rich metadata. In this paper, we intend to build a natural language based symbolic music generation system with our new public dataset MetaScore. MetaScore contains 963K musical scores paired with rich metadata collected from the MuseScore forum111https://musescore.com/ as well as extensive metadata such as genre, composer, complexity, time signature, key signature and user interaction statistics (e.g., number of views, likes and comments). In order to approach text to music generation, we further enhance the MetaScore dataset by completing missing genre metadata using a machine learning-based genre tagging algorithm and we leverage large language models to convert the metadata into natural language captions.
Model size | Public training data | Supports drums | Supports free-form text prompts | Controls | ||||
Instrument | Genre | Composer | Complexity | |||||
FIGARO [9] | 88.30M | ✓ | ✓ | |||||
MuseCoco [3] | 203M | ✓ | ✓ | ✓ | ✓ | |||
BART-based [10] | 139M | ✓ | ✓ | ✓ | ✓ | |||
MST-Tags | 87.36M | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |
MST-Text | 87.44M | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
Enabled by the metadata provided in MetaScore, we explore text-conditioned music generation that allows controls on instruments, genre, composer, and complexity with free-form text. With the LLM-enhanced dataset, we train a transformer-based text-to-music model using a pretrained large language model to encode the input text prompts. In addition, we train a transformer-based tags-to-music model by prepending the input tags to our proposed music representation. Leveraging the LLM-generated captions for training, the proposed text-to-music model achieves competitive performance against the tag-based model while offering a natural language-based interface that allows free-form text inputs. To evaluate our proposed models, we also compare them with an open-source text-to-symbolic music system [10]. In a listening study, we show that our proposed models outperform the baseline model in terms of coherence, arrangement, adherence, and overall quality. Our contributions can be summarized as follows:
We present a new publicly available dataset with musical scores paired with rich metadata and LLM-generated natural language captions.
We train two new models for tag- and text-based controllable symbolic music generation that support instrument, genre, composer and complexity controls.
The MetaScore dataset, including the musical scores, metadata and the LLM-generated captions, along with source code and pretrained models, will all be made publicly available upon acceptance. Audio samples can be found on our demo website.222https://wx83.github.io/MetaScore_Official/
II Prior Work
II-A Symbolic Music Datasets
We compare commonly used symbolic music datasets in WikiMusicText[13] pairs music with genre, composer, and captions. However, its publicly released version is small, and the musical scores are in ABC notation, which does not support multitrack music natively. Although MetaMIDI[15] comprises around 437K multitrack music pieces in MIDI format, it only includes genre and composer information, lacking natural language captions, which are significant for training text-to-music generation models. Although EMOPIA [11] provides emotion information, it is small and contains only pop music. In this work, we present a new and large multitrack and multi-genre symbolic music dataset with rich metadata, including genre, composer, complexity, key signature, time signature, user interaction statistics, free-form user annotated tags and pseudo captions.
II-B Controllable Symbolic Music Generation
Controllable symbolic music generation aims to generate symbolic music based on attribute values or free-form text descriptions. We compare our model with existing controllable music generation system in Table I. EMOPIA [11] is designed to generate music that aligns with specific emotional states, defined within the valence-arousal plane. This psychological model categorizes emotions by valence, indicating their positivity or negativity, and arousal, which measures their intensity from calm to excited. FIGARO [9] can generate samples based on a fine-grained description of the characteristics of the desired music. MuseCoco [3] extracts attributes from text descriptions to generate the desired music by training an attribute-to-music model. In this work, we explore free-form text conditioned music generation with LLM-generated natural language music captions. Additionally, we present a tag-conditioned music generation model that can generate music based on four conditions: genre, instrument, complexity and composer.
III MetaScore Dataset
In this section, we present the MetaScore dataset, a large collection of musical scores paired with rich metadata collected from an online music forum. Fig. 1 shows the statistics of important metadata available in MetaScore.
III-A Collecting and Preprocessing the Dataset
We collect 963K songs paired with musical scores and metadata from the MuseScore forum. We will refer to this original dataset as MetaScore-Raw. MetaScore-Raw contains extensive metadata such as genre, composer, complexity, key signature, time signature, tempo and user interaction statistics (e.g., number of views, likes and comments). We provide statistics of the metadata in Fig. 1, and we note that not all songs come with complete metadata. From the raw MSCZ files, we extract key signature, time signature, tempo and musical instrumentation. We only retain those instruments that are compatible with the General MIDI standard. Regarding composers, we first filter the composer tags to ensure they are formatted as human names and convert them to lowercase. We also standardize the names of well-known musicians to their full names; for instance, “mozart” is changed to “wolfgang amadeus mozart.”
III-B Inferring Missing Genre Tags in MetaScore-Raw
While MetaScore-Raw provides rich metadata information, we notice that not all songs come with complete metadata. For example, only 181K (18.8%) out of 963K songs in MetaScore-Raw contain genre metadata. As genre is one of the most intuitive ways for a user to control the style for a music generation system, we want to complete the genre information for songs without a genre label in MetaScore-Raw.
Therefore, we train a genre tagger that is based on the Multitrack Music Transformer (MMT) [1], where we remove the causal mask used for autoregressive modeling and append a multi-label classification layer. We select the threshold of the multi-label classification layer for each class based on the F1 score on the validation set. To evaluate the performance of genre tagging, we first compute the precision, recall and F1-score on the test set, where we achieve a micro-averaged precision of 61.94, recall of 63.03, and F1 score of 62.48. In addition, we conduct a subjective listening test to compare the quality of the auto-generated genre tags with the user-annotated tags in MetaScore-Raw. The 22 participants are instructed to answer the following question in a Likert scale of 1 to 5: “How well do you think this piece of music aligns with the following genre?”. From Table II, we can see that the auto-generated genre tags in MetaScore-Plus achieves a lower tags-music adherence compared to the ground truth tags in MetaScore-Genre, but the difference did not reach statistical significance in our setup.
III-C Generating Pseudo Captions using LLMs
To enable text-based downstream tasks (e.g., music captioning and text-to-music generation), we leverage large language models to convert the metadata into natural language captions. We follow LP-MusicCaps [6] and CLAP [5] and adopt an in-context learning-based approach [25] using a pretrained large language model. We form the input prompt string by combining genre, composer, copyright, complexity, and free-form user-specified tags. As shown in the demo page 2, we provide five examples of input-output pairs to facilitate in-context learning with Bloom, where the examples are used to provide guidance for the LLM to capture the one-to-many mapping between the input tags and natural language captions. We generate the pseudo captions using the Hugging Face API[31]. We exclude non-English and corrupted captions generated by Bloom and truncate the output sequence to a maximum of 32 tokens.
Type | Dataset | Adherence |
Ground truth genre tags | MetaScore-Genre | 3.11 0.49 |
Auto-generated genre tags | MetaScore-Plus | 3.05 0.54 |
LLM-generated captions | MetaScore-Plus | 3.23 0.49 |
We only include songs with auto-generated genre tags here.
To evaluate the quality of the generated pseudo captions, we also include this dataset in Table II and conduct a subjective listening test, where the participants are instructed to answer the following questions in a Likert scale of 1 to 5: “How well do you think this piece of music aligns with following text description?” As shown in Table II, the LLM-generated captions achieve the highest tags/text-music adherence across datasets, possibly because the participants prefer a natural language caption than a list of tags.
III-D Versions of MetaScore
We will release the following three versions of MetaScore:
MetaScore-Raw (963K): The raw MuseScore files and metadata scraped from the MuseScore forum.
Metascore-Genre (181K): A subset of MuseScore-Raw containing files with user-annotated genres. Additionally, we discard any songs composed by a composer that has less than 100 compositions in MetaScore-Raw. We also provide LLM-generated captions based on information extracted from the metadata in Metascore-Genre.
MetaScore-Plus (963K): MetaScore-Raw where missing genre tags are completed by the trained genre tagger described in Section III-B. We also provide LLM-generated captions based on information extracted from the metadata in MetaScore-Plus.
Due to copyright concerns, we will publicly release all music scores and metadata that are in the public domain (274K) or licensed with a Creative Commons licenses (52K). The rest of the dataset will be provided upon request for research purpose.
IV Methods
We represent a music piece as a one-dimensional array of integers using an event-based representation adapted from REMI+ [9] and MMT [1]. REMI+[9] represents notes with six consecutive tokens encoding note position, pitch, velocity, duration, instrument and time-signature information. However, it cannot provide control over tags such as genre, composer and complexity. MMT[1] represents a sequence of six-dimension events, with each event encoded as a tuple of variables (, , , , , ). However, MMT cannot model the interdependencies within these fields for a specific note as it predicts the six fields in parallel.
In this work, we adapt the REMI+ representation [9] to provide controls over genre, instrument, composer and complexity, while preserving the expressiveness offered by REMI+ [9]. Similar to MMT [1], we decompose note-on events to beat and position to reduce the size of the vocabulary and to help the model learn the rhythmic structure of music. In addition, we exclude the “tempo” and “chord” events as such information is generally unavailable in our dataset. Following REMI+ [9], we use beat, position, instrument, pitch and duration events for representing musical notes for non-drum tracks. We represent drum notes as beat, position, instrument, drum_pitch.
To enable free-form text controls, for each music piece with text, we use a pretrained sentence transformer [12] (specifically, the “all-MiniLM-L6-v2” version [30]) to extract the text embedding. Then we add a linear layer to project the text embedding to the input token embedding space, where the projected text embedding is added to the previous generated token embedding along with the positional encoding. Then we feed the encoded sequence into a decoder-only linear transformer. We will refer to this model as MetaScore Transformer-Text (MST-Text).
[…]
Dataset | Input type | Model size | Training samples | Coherence | Arrangement | Adherence | Overall quality | |
MST-Tags-Small | MetaScore-Genre | Tag | 87.36M | 150K | 3.87 0.36 | 3.98 0.38 | 3.86 0.38 | 3.57 0.37 |
MST-Tags | MetaScore-Plus | Tag | 87.36M | 901K | 4.01 0.37 | 4.06 0.39 | 3.60 0.49 | 3.66 0.45 |
BART-based [10] | Textune [10] | Text | 139M | 283K | 3.86 0.30 | 3.63 0.39 | 2.81 0.50 | 3.29 0.42 |
MST-Text | MetaScore-Plus | Text | 87.44M | 560K | 3.93 0.28 | 3.88 0.33 | 3.35 0.44 | 3.69 0.33 |
Pitch class entropy | Scale consistency | Groove consistency | |
MST-Tags-Small | 2.88 0.08 | 0.89 0.02 | 0.92 0.01 |
MST-Tags | 2.93 0.07 | 0.89 0.02 | 0.90 0.01 |
MST-Text | 2.70 0.06 | 0.95 0.01 | 0.92 0.01 |
Ground truth | 2.67 0.06 | 0.95 0.01 | 0.92 0.01 |
V Experiments and Results
V-A Baseline Models
We compare our method to an existing work for symbolic music with ABC notations with the same prompts that we randomly pick from our test set. BART-base [10] investigate the efficacy of pre-trained checkpoints in text-to-music generation. We generate music with BART-base model in ABC notation using its Hugging Face API [32]. For a fair comparison, we convert the generated ABC notation to MIDI using Melobytes, an online software which supports multitrack conversion [33]. We intended to compare our proposed model to MuseCoco[3]. However, not all free-form natural language captions in our test set can conform to their template.
V-B Objective Evaluation
Following [1, 27, 26], we use pitch class entropy, scale consistency and groove consistency to evaluate generated music. We randomly pick 100 samples from the test set to compute these metrics, and we consider a closer value to that of the ground truth better. We report in Table IV the pitch class entropy, scale consistency and groove consistency of MST-Tags-Small, MST-Tags, MST-Text and the ground truth. We consider the generated music better if the corresponding value is closer to that of the ground truth. We observe that MST-Tags-Small and MST-Tags lead to values significantly different from those of the ground truth in terms of pitch class entropy and scale consistency.
Further, MST-Text is able to generate music that is statistically close to the ground truth in terms of all three metrics. However, we note that none of these objective evaluation metrics measures the tags/text-to-music adherence, and we will examine this in the subjective listening test.
V-C Subjective Evaluation
We conduct a subjective test where 22 participants are instructed to evaluate five songs under each scenario. Out of the 22 participants, 19 people have experience in playing instruments, with two being professional musicians. We ask the participants to evaluate the audio samples in terms of coherence, arrangement, adherence and overall quality in a Likert scale of 1 to 5.
We report the subjective evaluation results in Table III. When contrasting MST-Tags-Small with MST-Tags, we observe that MST-Tags achieves better performance in coherence and arrangement, but we see a decrease in adherence, possibly due to the incorporation of some auto-generated tags. This comparison illustrates the trade-off between employing a smaller, high-quality dataset (MetaScore-Genre) versus a larger yet noisy dataset (MetaScore-Plus). However, comparing the overall quality score of MST-Tags-Small and MST-Tags, we see that training with a larger dataset leads to an increase in the overall quality of music generation.
For text-conditioned music generation, MST-Text outperforms the BART-based[10] approach, in terms of coherence, arrangement, adherence, and overall quality. We observe that text-conditioned system MST-Text has a lower adherence against MST-Tags-Small and MST-Tags. However, Table II shows our generated captions keep all the information with in-context learning. This implies that text-to-music generation is a more challenging task than tag-to-music generation because additional effort is required to interpret free-form text for music generation.
Overall, the tag-conditioned music generation systems, MST-Tags and MST-Tags-Small, perform very well, suggesting the high quality of our constructed dataset. In addition, MST-Text achieves the highest overall quality, which shows that our text-conditioned system performs similarly well with our tag-conditioned system. This showcases the effectiveness of our approach that leverages a large language model to generate natural language captions for end-to-end training of text-to-music generation models.
VI Conclusion
In this work, we have presented a new music generation model that learns to generate symbolic music from free-form text, allowing controls over instruments, genre, composer, complexity and other free-form music descriptors. In addition, the LLM-generated pseudo captions contain information provided in free-form user-annotated tags, which can poses a challenge to systems that adopt a predefined set of tags, including our tag-to-music model and the two-stage approach proposed in [3]. Our objective and subjective evaluation results show the effectiveness of the proposed tags-to-music and text-to-music models, both outperforming a baseline text-to-music model [10]. In addition, the proposed text-to-symbolic music generation model trained with LLM-generated pseudo captions achieves competitive performance against the proposed tags-to-music model trained using only the ground truth tags.
Finally, we would like to point out that, as a generative model trained on copyrighted material, our proposed model has the potential to generate samples that could lead to copyright infringement. However, when used properly and with cautions, a text-to-music generation system can also make a positive impact to the society by enabling new opportunities and interfaces for music creation, as demonstrated in [28]. Given the editable nature of symbolic music, we hope our proposed text-to-symbolic music models to open up new pathways towards human-AI music co-creation.